how many gallons of water that is 74% of alcohol must be mixed with gallons of one that is 90% alcohol in order to obtain a mixture that is 84% alcohol?
Answer by stanbon(75887)

how many gallons of water that is 74% of alcohol must be mixed with gallons of one that is 90% alcohol in order to obtain a mixture that is 84% alcohol?
-------------------
That depends on how many gallons of the mixture you want.
Let's say you want 100 gallons.
Equation:
alcohol + alcohol = alcohol
0.74x + 0.90(100-x) = 0.84*100
---
Multiply thru by 100 to get:
74x + 9000-90x = 8400
-16x = -600
x = 37.5 gallons (amt of 74% solution in the 84% mixture)
100-x = 62.5 gallons (amt. of 90% solution in the 84% mixture)
================
Cheers,
Stan H.